A radioisotope decays to give an alpha particle and protactinium-233. What was the original nucleus


1
Expert's answer
2021-12-09T10:09:02-0500

Solution:

The original nucleus is Neptunium- 237 , 

Explanation

Let the original Nucleus be X.

We are provided with the information that after thee decay of X, we get an alpha particle and a protactinium-233 nucleus ,

We know that for Alpha particle, Mass number A= 4 and Atomic number Z= 2

And for protactinium- 233, Mass number A= 233 and atomic number Z= 91

Hence the decay of X can be represented as follows

As we can see the above nuclear reaction is an example of alpha decay process.

So here the Atomic number Z of the original nucleus X will be

Z= 91 + 2=93

And the mass number A of the original nucleus X will be

A= 233+ 4=237

So, Element X can be represented in the isotopic notation as 

The atomic number and mass number actually corresponds to that of Neptunium . Hence, nucleus X is 

Hence, the original nucleus would be Neptunium-237,